OpenStudy (anonymous):

7x+2y=5 can anyone show me how to solve for x?

OpenStudy (anonymous):

Don't post the same question multiple times. And try and explain more. This is a pretty equation with no question.

OpenStudy (anonymous):

it's no answer but x = (5 - 2y)/7

OpenStudy (anonymous):

u get the other term on Left Hand Side (LHS) that is 2y on the RHS which becomes 5-2y so 7x=5-2y then divide by 7 from both sides to get x on its own and then u will have solved for x.

OpenStudy (anonymous):

7x+2y=5 put x=1 & y=-1 then get o/p=5

OpenStudy (anonymous):

i think to find the value of x just simply substitute 0 in the value of y . so its x=5/7.